### v1

@@ Appendix A Example of unitary and causal amplitude violating Froissart-Martin bound
In Fig. 5 we showed the behavior of $\sqrt{s}\Delta T(s)$ for the amplitude in (39). We found that indeed macroscopic causality is satisfied everywhere from zero to infinite energies since $|\sqrt{s}\Delta T(s)|<1$ for any $s$.
Figure 5: Plot of $\sqrt{s}\Delta T$ for the amplitude (39) with $\gamma=0.1$. This quantity is bounded from below and from above, and never becomes smaller than $-1$. This is consistent with the causality condition that forbids resolvable time advances. However, in this case, asymptotic causality is not satisfied.
Appendix B Derivation of the analytic bound on $g_{3}$
In this appendix, we provide more details about the method that was used to derive the bounds involving the coupling $g_{3}$. We refer to [101, 69, 102] for more detailed derivation and explanations.
@@ Appendix B Derivation of the analytic bound on $g_{3}$
$\left|\begin{matrix}\int f_{1}^{2}&\int f_{1}f_{2}&\int f_{1}f_{3}\\[4.2679pt] \int f_{2}f_{1}&\int f_{2}^{2}&\int f_{2}f_{3}\\[4.2679pt] \int f_{3}f_{1}&\int f_{3}f_{2}&\int f_{3}^{2}\end{matrix}\right|\geq 0\quad\Rightarrow\quad\left|\begin{matrix}a_{2,0}&a_{3,0}&a_{3,1}\\[4.2679pt] a_{3,0}&a_{4,0}&a_{4,1}\\[4.2679pt] a_{3,1}&a_{4,1}&a_{4,2}\end{matrix}\right|\geq 0\,,$ (47)
@@
$\left|\begin{matrix}a_{2,0}&a_{3,0}\\[4.2679pt] a_{3,0}&a_{4,0}\\ \end{matrix}\right|\geq 0\,,\qquad\left|\begin{matrix}a_{2,0}&a_{3,1}\\[4.2679pt] a_{3,1}&a_{4,2}\end{matrix}\right|\geq 0\,,\qquad\left|\begin{matrix}a_{4,0}&a_{4,1}\\[4.2679pt] a_{4,1}&a_{4,2}\end{matrix}\right|\geq 0\,.$ (48)
The final condition follows from the fact that $L^{2}\in\{0,6,20,42,\dots\}$ is discrete-valued, thus we have $(L^{2}-6)(L^{2}-20)\geq 0$ for all possible values of $L^{2}$ in the discrete set. Thus, we have
$\left[\frac{(L^{2}-6)(L^{2}-20)}{\mu^{5}}\right]\geq 0,\qquad a_{4,2}-26a_{4,1}+120a_{4,0}\geq 0.$ (49)
@@
If we define the function $f(\tilde{a}_{3,0})\equiv\frac{20}{3}\tilde{a}_{3,0}+\sqrt{\frac{80}{9}(\tilde{a}_{4,0}-\tilde{a}_{3,0}^{2})}$, then the most loose bound corresponds the maximum of $f(\tilde{a}_{3,0})$. This maximum value is given by $\tilde{a}_{3,0}=\sqrt{\frac{5}{6}\tilde{a}_{4,0}}$. In addition, from the other constraints in Eqs. (46) and (48), we also have $\tilde{a}_{4,0}\leq\tilde{a}_{3,0}\leq\sqrt{\tilde{a}_{4,0}}$, which means that $f(\tilde{a}_{3,0})$ can also reach maximum at the boundary of the allowed range for $\tilde{a}_{3,0}$. Therefore, we need to consider these two regions separately,
$0<\tilde{a}_{3,1}<\sqrt{\frac{160}{3}\tilde{a}_{4,0}}$ $\text{for}\qquad 0<\tilde{a}_{4,0}\leq\frac{5}{6},$ (53)
$0<\tilde{a}_{3,1}<\sqrt{\frac{80}{9}(\tilde{a}_{4,0}-\tilde{a}_{4,0}^{2})}+\frac{20}{3}\tilde{a}_{4,0}$ $\text{for}\qquad\frac{5}{6}<\tilde{a}_{4,0}\leq 1.$
The shape carved by these inequalities is called a-geometry, the result is general, and does not rely on the definition of the brackets in Eq. (45). In our study that involves non-local amplitudes, we need to replace $\tilde{a}_{4,0}=2\tilde{g}_{4}+\tilde{\alpha}$ and $\tilde{a}_{3,1}=-\frac{1}{2}\tilde{g}_{3}+\frac{3}{2}\tilde{a}_{3,0}$ into Eq. (51) and then project out $\tilde{a}_{3,0}$ to get the result in Eq. (29).

### v2

@@
We found it much easier to construct examples of amplitudes that satisfy only the macroscopic causality condition. In fact, it suffices to require them to be smooth bounded functions for real and positive values of $s$. In this sense, the unitarity condition implies macroscopic causality, while the asymptotic causality seems to be a much stronger requirement than a combination of analyticity and partial-wave unitarity.
Derivation of the analytic bound on $g_{3}$
We provide here more details about the method that was used to derive the bounds involving the coupling $g_{3}$. We refer to [98, 69, 99] for more detailed derivation and explanations. To derive a constraint on the coupling $\tilde{g}_{3}$, we also need the partial-wave decomposition for the discontinuities in $d=4$ spacetime dimensions, i.e.
where $n_{l}=16\pi(2l+1),$ $f_{l}(s)$ is the partial-wave amplitude, $\rho_{l}(s)={\rm Im}f_{l}(s)$ is the spectral density, $\mathcal{P}_{l}(x)$ are the Legendre polynomials and the angular momentum $l$ only takes even values because we are working with a scalar field theory.
If we define
take derivatives with respect to $s$ and $t$, and set $s\rightarrow 0$, $t\rightarrow 0$ afterwards, we can write
$\begin{split}&g_{2}=\left[\frac{1}{\mu^{3}}\right]\,,\,\,\,g_{3}=\left[\frac{3}{\mu^{4}}\right]-\left[\frac{2L^{2}}{\mu^{4}}\right]\,,\,\,\,\\ &g_{4}=\left[\frac{1}{2\mu^{5}}\right]-\left[\frac{\alpha}{2\mu^{3}}\right].\end{split}$ (40)
where $L^{2}\equiv l(l+1)$. Crossing symmetry implies that the relation for the coupling $g_{4}$ in the EFT expansion can be derived in two different ways. The first way is differentiating four times with respect to $s$ and then taking $s,t\rightarrow 0$. The second way is differentiating twice with respect to $s$ and twice with respect to $t$ and then taking $s,t\rightarrow 0$. Since the two procedures must be equivalent, we get the so-called null constraint:
$\left[\frac{L^{2}(L^{2}-8)}{\mu^{5}}\right]=0\,.$ (41)
@@
$\left|\begin{matrix}\int f_{1}^{2}&\int f_{1}f_{2}&\int f_{1}f_{3}\\[4.2679pt] \int f_{2}f_{1}&\int f_{2}^{2}&\int f_{2}f_{3}\\[4.2679pt] \int f_{3}f_{1}&\int f_{3}f_{2}&\int f_{3}^{2}\end{matrix}\right|\geq 0\,\Rightarrow\,\left|\begin{matrix}a_{2,0}&a_{3,0}&a_{3,1}\\[4.2679pt] a_{3,0}&a_{4,0}&a_{4,1}\\[4.2679pt] a_{3,1}&a_{4,1}&a_{4,2}\end{matrix}\right|\geq 0\,,$ (44)
@@
$\left|\begin{matrix}a_{2,0}&a_{3,0}\\[4.2679pt] a_{3,0}&a_{4,0}\\ \end{matrix}\right|\geq 0\,,~{}\left|\begin{matrix}a_{2,0}&a_{3,1}\\[4.2679pt] a_{3,1}&a_{4,2}\end{matrix}\right|\geq 0\,,~{}\left|\begin{matrix}a_{4,0}&a_{4,1}\\[4.2679pt] a_{4,1}&a_{4,2}\end{matrix}\right|\geq 0\,.$ (45)
The final condition follows from the fact that $L^{2}\in\{0,6,20,42,\dots\}$ is discrete-valued, thus we have
$(L^{2}-6)(L^{2}-20)\geq 0$ (46)
for all possible values of $L^{2}$ in the discrete set. Thus, we have
$\begin{split}&\left[\frac{(L^{2}-6)(L^{2}-20)}{\mu^{5}}\right]\geq 0,~{}\\ &a_{4,2}-26a_{4,1}+120a_{4,0}\geq 0.\end{split}$ (47)
@@
If we define the function
$f(\tilde{a}_{3,0})\equiv\frac{20}{3}\tilde{a}_{3,0}+\sqrt{\frac{80}{9}(\tilde{a}_{4,0}-\tilde{a}_{3,0}^{2})},$ (51)
then the most loose bound corresponds the maximum of $f(\tilde{a}_{3,0})$. This maximum value is given by $\tilde{a}_{3,0}=\sqrt{\frac{5}{6}\tilde{a}_{4,0}}$. In addition, from the other constraints in Eqs. (43) and (45), we also have $\tilde{a}_{4,0}\leq\tilde{a}_{3,0}\leq\sqrt{\tilde{a}_{4,0}}$, which means that $f(\tilde{a}_{3,0})$ can also reach maximum at the boundary of the allowed range for $\tilde{a}_{3,0}$. Therefore, we need to consider these two regions separately,
$0<\tilde{a}_{3,1}<\sqrt{\frac{160}{3}\tilde{a}_{4,0}},\qquad\qquad\qquad 0<\tilde{a}_{4,0}\leq\frac{5}{6},$ (52)
$\tilde{a}_{3,1}<\sqrt{\frac{80}{9}(\tilde{a}_{4,0}-\tilde{a}_{4,0}^{2})}+\frac{20}{3}\tilde{a}_{4,0},~{}~{}\frac{5}{6}<\tilde{a}_{4,0}\leq 1.$
The shape carved by these inequalities is called a-geometry, the result is general and does not rely on the definition of the brackets in Eq. (42). In our study that involves non-local amplitudes, we need to replace $\tilde{a}_{4,0}=2\tilde{g}_{4}+\tilde{\alpha}$ and $\tilde{a}_{3,1}=-\frac{1}{2}\tilde{g}_{3}+\frac{3}{2}\tilde{a}_{3,0}$ into Eq. (49) and then project out $\tilde{a}_{3,0}$ in order to get our final result for the bound on $\tilde{g}_{3}$. It is worth mentioning that adding higher-order null constraints would only lead to a very small improvement of the bound in its upper left corner. This does not affect any of our conclusions.